probability a student will pass an economics exam is 1/2
probability the student pass economics and management exam is 3/5
probability that the student pass economics or management exam is 4/5
find the probability the students pass a management exam

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

P(A)+P(B)-P(A+B)=P(A or B)
P(E)=0.5
0.5+P(M)-P(E+M)=P(E or M)
P(E+M)=0.6
P(E or M)=0.8
so
0.5+P(M)-0.6=0.8
P(M)=0.9

Answer by ikleyn(52809)

.
probability a student will pass an economics exam is 1/2
probability the student pass economics and management exam is 3/5
probability that the student pass economics or management exam is 4/5
find the probability the students pass a management exam
~~~~~~~~~~~~~~~~~


The problem is posed  FATALLY  INCORRECTLY.


The probability of the intersection of two events  ( passing  Economics  AND  Management,  3/5 )   CAN  NOT  BE  GREATER

the probability of any of the two basic events  ( passing  Economics,  1/2 ).


Posed this way as written in the post,  the  "problem"  is a  FAKE  and is good for re-cycling,   ONLY.